7)Find and compare the future value after two years of a deposit of $100 attracting interest at a rate a) annually and b) semiannually.
8)Which will deliver a higher future value after one year, a deposit of $1, 000 attracting interest at 15% compounded daily, or at 15.5% compounded semi-annually?
9)What initial investment subject to annual compounding at 12% is needed to produce $1, 000 after two years?
10)Which will deliver a higher future value after one year, a deposit of $1, 000 attracting interest at 15% compounded daily, or at 15.5% compounded semi-annually?
11)What initial investment subject to annual compounding at 12% is needed to produce $1, 000 after two years?
12)Find the present value of $100, 000 to be received after 100 years if the interest rate is assumed to be 5% throughout the whole period and a) daily or b) annual compounding applies.
1
Expert's answer
2020-06-10T20:00:54-0400

7)

"FV=100(1+0.1)^2=121"


"FV=100(1+\frac{0.1}{2})^4=121.55"


once every six months more

8)

"FV=1000(1+\frac{0.15}{365})^{365}=1161.79"

"FV=1000(1+\frac{0.155}{2})^{2}=1 161.00"


accrued daily will give more


9)

"1000=PV(1+0.12)^2=797.19"


10)


"FV=1000(1+\frac{0.15}{365})^{365}=1161.79"


"FV=1000(1+\frac{0.155}{2})^{2}=1 161.00"

accrued daily will give more


11)

"1000=PV(1+0.12)^2=797.19"

12)

a)

"100 000=PV(1+\frac{0.05}{365})^{36500}=673.93"

b)"100 000=PV(1+{0.05}^{100}=760.45"